# Latent BKI: Open-Dictionary Continuous Mapping in Visual-Language Latent Spaces with Quantifiable Uncertainty

### v1

@@ Abstract
This paper introduces a novel probabilistic mapping algorithm, Latent BKI, which enables open-vocabulary mapping with quantifiable uncertainty. Traditionally, semantic mapping algorithms focus on a fixed set of semantic categories which limits their applicability for complex robotic tasks. Vision-Language (VL) models have recently emerged as a technique to jointly model language and visual features in a latent space, enabling semantic recognition beyond a predefined, fixed set of semantic classes. Latent BKI recurrently incorporates neural embeddings from VL models into a voxel map with quantifiable uncertainty, leveraging the spatial correlations of nearby observations through Bayesian Kernel Inference (BKI). Latent BKI is evaluated against similar explicit semantic mapping and VL mapping frameworks on the popular MatterPort-3D and Semantic KITTI data sets, demonstrating that Latent BKI maintains the probabilistic benefits of continuous mapping with the additional benefit of open-dictionary queries. Real-world experiments demonstrate applicability to challenging indoor environments.
@@ I Introduction
Robots require informative world models to autonomously navigate the world, commonly known as maps. Mapping methods represent the geometry of the robot’s surroundings and often include semantic information relevant to robotic task success. While some works have proposed mapless autonomous navigation [1, 2], maps are commonly used in robotics due to the ability to leverage temporal information.
Maps are also capable of storing a high level of scene understanding with multi-modal information such as occupancy, semantics, traversability, and uncertainty. Early mapping algorithms created probabilistic 2D occupancy grids by extracting free space samples through ray tracing [3]. As deep neural networks have rapidly progressed, mapping algorithms expanded to store logistic regression outputs from neural networks in 3D semantic maps. However, features in the latent space of neural networks store rich context, which can be decoded as useful information for other tasks. Additionally, real-world environments contain complex and detailed scenes that cannot be captured through closed-dictionary maps.
@@
Recently, deep learning has produced foundation models trained on large, varied data sets with the reported ability to generalize to out-of-distribution data, solving a limitation of previous semantic segmentation neural networks [4, 5, 6]. Additionally, such open-dictionary segmentation networks can share a feature space with large language models [7], creating a class of networks known as Vison-Language (VL) models [8]. While several methods for VL mapping have been proposed, classical robotic challenges of continuous mapping and quantifying uncertainty remain.
Quantifying uncertainty is critical for robotics applications, as observations of the world are limited and noisy, both from sensor noise as well as neural network errors. Probabilistic mapping seeks to quantify uncertainty in the world model by tracking the number of observations of the environment, as well as the variance or stochasticity of the observations. Additionally, in practice, observations of the environment may be sparse, leading to incomplete maps.
Completing maps from sparse data is known as the task of scene completion, which seeks to leverage exteroceptive information to form a more complete scene. While one branch of scene completion seeks to complete maps implicitly through neural networks [9, 10], a probabilistic approach to scene completion is known as continuous mapping, which leverages spatial observations to complete representations of the environment [11]. One benefit of continuous mapping is quantifiable uncertainty, as the influence of points in the map is modeled through spatial relationships. However, continuous semantic mapping has been limited to closed-dictionary categorical spaces, which is impractical for complex environments.
This paper proposes an extension of continuous mapping to the latent space of neural networks, enabling open-dictionary continuous mapping. The proposed method, LatentBKI, performs Bayesian Kernel Inference (BKI) [12] over the latent embedding space to probabilistically complete scenes from sparse data with quantifiable uncertainty through conjugate priors. We compare our method against the pioneering work VL-Map, which introduced a method for mapping in a latent space from VL networks but does not quantify uncertainty or leverage spatial relations. We also compare our method against closed-dictionary continuous mapping methods, demonstrating that our method enables open-dictionary mapping without losing performance and with similar abilities to quantify uncertainty. Finally, we evaluate our method on real-world indoor scenes, highlighting the advantage of leveraging open-dictionary segmentation networks for continuous mapping. To summarize, our contributions are:
@@ II-A Continuous Semantic Mapping
Robots require advanced levels of scene understanding to plan, including knowledge of the geometry and semantic labels of objects and uncertainty associated with the objects to avoid failure due to mistaken object identity. Often, these approaches use task-dependent object designations as semantic labels [13, 14] such as abstract topological information [15] or material classifications [16]. State-of-the-art methods for semantic classification almost exclusively rely on neural networks to produce pixel-wise semantic predictions [17, 18, 19]. Neural networks can successfully categorize objects. Robotics research has focused on incorporating segmentation predictions into maps via semantic label fusion [20, 21]. Recent methods aim to quantify uncertainty through Bayesian inference [22, 23] by iteratively fusing semantic estimates projected onto a geometric map.
Kernel-based inference schemes have had notable success [24, 25] in probabilistic semantic mapping. Bayesian Kernel Inference (BKI), proposed by [12], approximates the spatial influence of points at model selection through the usage of a kernel. BKI is an approximation of Gaussian Processes, which are effective at continuous mapping yet suffer from a cubic computational complexity [11, 26, 27]. Effectively, the kernel defines the shape or distribution of a point, deemed the extended likelihood, and can be applied to create an efficient, closed-form Bayesian update with more complete maps and quantifiable uncertainty [28]. While BKI has been applied effectively to semantic mapping [29, 25, 30], semantic maps are inherently limited to a closed set of pre-specified categories. In contrast, we propose to extend the literature of continuous mapping to the latent space of neural networks, allowing for open-vocabulary inference within the latent space of VL models.
Figure 2: This demonstrates the overall pipeline of LatentBKI. (a) The input to LatentBKI is 3D points, which can be from LiDAR, RGB-D, or any exteroceptive sensor with 3D input. (b) Points are then processed by an off-the-shelf neural network, which encodes each point into a latent space. (c) By adopting a Gaussian likelihood over the point-wise features, we perform closed-form Bayesian inference on a voxel map where each voxel contains parameters modeling the conjugate prior of the multivariate Gaussian distribution. Additionally, instead of only considering points which fall within a voxel, we consider nearby points weighted through a kernel function. (d) The posterior predictive distribution of each voxel in latent space can then be decoded using the decoder of the neural network, enabling the computation of open-dictionary segmentation predictions with expectation and uncertainty.
@@ II-B Vision Language Mapping
Rapidly improving Large Language Models (LLMs) demonstrating remarkable generalizable capabilities have motivated the advent of vision-language models (VLM) with shared latent space for both images and texts [8]. The pioneering method CLIP successfully represents visual and textual information in the same embedding dimension through contrastive learning [6]. Trained on a large dataset of image-text pairs, CLIP learns embed features from visual or textual information in a shared feature space, where similarity is measured by a cosine similarity function [6]. Several following works have improved upon the data efficiency of CLIP [31] and image contextualization [32].
Based on the success of VLMs and their great zero-shot performance, recent robotics research has focused on open-dictionary mapping which operates in the latent space of VLMs and can create segmentation predictions from language descriptors [7]. Approaches like LM-Nav [33], CoW [34], NLMap [35] have fused VLMs to enable robots to understand and navigate new environments. One of the pioneering approaches is VLMaps [36], which showed outstanding performance on zero-shot navigation tasks. It uses the language-driven semantic segmentation network LSeg [7] to generate feature embeddings for pixels in images the robot observes. Pixel-wise latent embeddings are then stored in a 3D voxel map through a weighted averaging scheme, where weights are dependent on pixel depths and points at extreme distances are excluded [37, 38]. The latent expectation of each voxel can then be decoded into per-category scores given the language embeddings of a set of categories, thereby enabling open-dictionary queries. While successful, this approach loses the ability to quantify uncertainty and fill in gaps in the map from probabilistic continuous mapping, as discussed previously. Therefore, we propose to extend closed-dictionary continuous semantic mapping to open-dictionary VL maps.
@@ III-A Latent Mapping Representation
Our map representation consists of a voxel map with voxels * located at position $x_{*}$. At each time-step our map is provided a set of points $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{3}$ is the position of point $i$ and $y_{i}\in\mathbb{R}^{C}$ is the corresponding latent feature of point $i$. From these points, our goal is to probabilistically update the latent parameters of each voxel, $y_{*}$, to obtain the posterior.
In order to accomplish this goal, we first define a Gaussian likelihood over the feature space, such that:
$p(y_{i}|\mu_{i},\Sigma_{i})=\mathcal{N}(y_{i};\mu_{i},\Sigma_{i}).$ (3)
@@
Since the features originate from a neural network, the likelihood defines an unknown expectation and variance which the point’s features are sampled from. Similarly, we can define the points observed within a voxel * according to the same likelihood distribution. From the likelihood, we can write an expression for the posterior of the latent parameters $\theta_{*}=\{\mu_{*},\Sigma_{*}\}$ of voxel * using Bayes’ rule as:
$p(\theta_{*}\mid x_{*},\mathcal{D})\propto p(\mathcal{D}\mid\theta_{*},x_{*})p(\theta_{*}\mid x_{*}).$ (4)
@@
In order to model the distribution over the parameters $\theta_{*}$ of the voxel, which themselves define a multivariate Gaussian distribution, we adopt the conjugate prior of the multivariate Gaussian distribution, the normal-inverse Wishart distribution. The normal-inverse Wishart distribution defines a distribution over the multivariate Gaussian with unknown mean $\mu_{*}$ and covariance $\Sigma_{*}$ through parameters $\mu^{\prime}_{*},\Psi_{*},\lambda_{*},v_{*}$ as:
@@
$p(\mu_{*},\Sigma_{*})=\mathcal{N}\left(\mu_{*};\mu^{\prime}_{*},\frac{\Sigma_{*}}{\lambda_{*}}\right)\cdot\mathcal{W}^{-1}(\Sigma;\Psi_{*},v_{*}),$ (5)
@@
where the hyper-parameters represent the prior expectation of the mean ($\mu^{\prime}_{*}$), the expectation of the covariance ($\Psi_{*}$), and the confidence in the mean and covariance estimates ($\lambda_{*},v_{*}$). In our case, $\lambda_{*}$ and $v_{*}$ are equal and correspond to weighted counts of the total observations.
@@
Although the conjugate prior provides a closed-form solution for updating the multivariate normal distribution parameters for each voxel, it does not consider the spatial locations of points. Intuitively, points that are closer to the centroid of the voxel should have a higher influence, while points that are further from the voxel centroid should have a lower influence. Additionally, only updating the voxels which points fall into can lead to sparse maps, as previously noted. Therefore, we adopt the solution of [12], which defines an extended likelihood distribution that considers the spatial relationship of points to voxels through the use of a kernel function $k(x_{i},x_{*})$ as:
$p(y_{i}|x_{i},\theta_{*},x_{*})\propto p(y_{i}|\theta_{*})^{k(x_{i},x_{*})}.$ (6)
@@
The only requirements when defining the extended likelihood are that
$k(x,x)=1\,\forall\,x\quad\text{and}\quad k(x,x_{*})\in[0,1]\,\forall\,(x,x_{*}).$ (7)
Applied to the previously defined Gaussian likelihood, the extended likelihood can be written as:
$p(y_{i}|\theta_{*},x_{i},x_{*})=\mathcal{N}\left(y_{i};\mu_{*},\frac{\Sigma_{*}}{k(x_{i},x_{*})}\right).$ (8)
Effectively, observations close to the voxel centroid have lower variance than observations of further points, resulting in more influence or gain. Following Semantic BKI, we use a symmetric sparse kernel [39] as the kernel function for direct comparison. Substituting the extended likelihood into (4), we can now define a spatial update over the voxel parameters as:
$p(\theta_{*}|x_{*},y_{1:N},x_{1:N})\propto\left[\prod_{i=1}^{N}p(y_{i}|\theta_{*},x_{i},x_{*})\right]p(\theta_{*}|x_{*}).$ (9)
@@ III-B Latent Mapping Update
First, we initialize the confidence over the mean and covariance of each voxel to a non-informative value of $\lambda_{*}\approx 0$. As points are observed, the value of $\lambda_{*}$ increases, indicating more confidence in the expected mean and covariance of the voxel. At time-step $t$, voxels are parameterized by prior $\mu_{*}^{t-1}$, $\Psi_{*}^{t-1}$ and confidence $\lambda_{*}^{t-1}$, with input points $\mathcal{D}$.
The influence of the new observations is calculated by:
$\bar{k}_{*}=\sum_{i=1}^{N}k(x_{*},x_{i}),$ (10)
where the kernel function measures the influence of point $i$ over voxel *. The confidence in the mean and covariance is updated:
$\lambda_{*}^{t}=\lambda_{*}^{t-1}+\bar{k}_{*}.$ (11)
Input observations are then used to compute the new mean as a running average:
$\bar{y}_{*}=\sum_{i=1}^{N}\frac{k(x_{*},x_{i})}{\bar{k}_{*}}y_{i},\quad\mu_{*}^{t}=\frac{\lambda_{*}^{t-1}\mu_{*}^{t-1}+\bar{y}_{*}\bar{k}_{*}}{\lambda_{*}^{t}}.$ (12)
@@
As new points are obtained, the above process is repeated to update the mean, covariance, and confidence level. For simplicity, we assume a diagonal covariance, which significantly reduces the number of parameters stored per voxel. Diagonal covariances are common in the feature space and are used in variational auto-encoders (VAEs).

### v2

@@ Abstract
This paper introduces a novel probabilistic mapping algorithm, LatentBKI, which enables open-vocabulary mapping with quantifiable uncertainty. Traditionally, semantic mapping algorithms focus on a fixed set of semantic categories which limits their applicability for complex robotic tasks. Vision-Language (VL) models have recently emerged as a technique to jointly model language and visual features in a latent space, enabling semantic recognition beyond a predefined, fixed set of semantic classes. LatentBKI recurrently incorporates neural embeddings from VL models into a voxel map with quantifiable uncertainty, leveraging the spatial correlations of nearby observations through Bayesian Kernel Inference (BKI). LatentBKI is evaluated against similar explicit semantic mapping and VL mapping frameworks on the popular Matterport3D and Semantic KITTI data sets, demonstrating that LatentBKI maintains the probabilistic benefits of continuous mapping with the additional benefit of open-dictionary queries. Real-world experiments demonstrate applicability to challenging indoor environments.
Index Terms:
Mapping, Semantic Scene Understanding, Deep Learning for Visual Perception.
@@ I Introduction
Robots require informative world models to autonomously navigate the world, commonly known as maps. Mapping methods represent the geometry of the robot’s surroundings and often include semantic information relevant to robotic task success. While some works have proposed mapless autonomous navigation [1, 2], maps are commonly used in robotics due to the ability to leverage temporal information within an interpretable world model.
Maps are also capable of storing a high level of scene understanding with multi-modal information such as occupancy, semantics, traversability, and uncertainty. As deep neural networks have rapidly progressed, mapping algorithms have also evolved from binary occupancy grids [3] to model higher levels of scene understanding such as through semantic information [4]. However, real-world environments contain complex and detailed scenes that cannot be captured through closed-dictionary semantic maps.
@@
Recently, deep learning has produced foundation models trained on large, varied data sets with the reported ability to generalize to out-of-distribution data, solving a limitation of previous semantic segmentation neural networks [5, 6]. Of particular interest to the robotics community, Vision-Language (VL) models [7] extend vision processing to share a feature space with large language models, enabling applications such as open-dictionary semantic segmentation networks [8]. Following the success of open-dictionary segmentation networks, several robotic mapping papers have proposed to integrate open-dictionary segmentation features within a robotic map in order to enable language-based queries and navigation [9, 10, 11, 12, 13]. Within robotic open-dictionary mapping networks, a common approach is to fuse input points into the map through a simple moving average scheme, which does not require additional training [10, 11, 13]. Inspired by the success of the simple averaging technique, we propose to leverage the structure of continuous mapping to enable spatial smoothing and uncertainty quantification.
Continuous mapping is an approach to probabilistic mapping which leverages Bayesian inference and spatial context to create complete maps with uncertainty [14, 15, 16]. Quantifiable uncertainty is critical for robotics applications, as observations of the world are often limited and noisy, leading to errors and incomplete maps. Through uncertainty quantification in the form of variance, planning algorithms can identify potentially dangerous conditions [17] and the optimal trajectories to obtain new measurements [18, 19]. Additionally, in the real world data is often sparse whether from sensors or views, leading to incomplete map representations. An efficient approach to continuous mapping known as Bayesian kernel inference (BKI) leverages spatial context to update nearby unobserved voxels through the use of an extended likelihood function defined by a kernel [20]. BKI yields an efficient probabilistic update, however has not been studied as a possible solution for open-dictionary mapping.
In this paper we propose an extension of continuous mapping to the latent space of neural networks called LatentBKI, which enables open-dictionary continuous mapping with spatial smoothing and quantifiable uncertainty. Compared to prior works in open-dictionary mapping, our method also leverages a weighted average to calculate the per-voxel expectation, however enables uncertainty quantification and fills in gaps within the map through BKI [20]. Compared to prior research in continuous mapping for robotic maps, LatentBKI proposes a Gaussian likelihood conjugate pair for latent space measurements, enabling open-dictionary mapping and inference without any loss in performance.
We evaluate our method against VLMap, which leverages a moving average recursive approach and is directly comparable to the expectation produced by LatentBKI. We also quantitatively evaluate our method against closed-dictionary continuous mapping methods, demonstrating that LatentBKI enables open-dictionary mapping without losing performance and while maintaining the ability to quantify uncertainty. Finally, we evaluate our method on real-world indoor scenes, highlighting the advantage of leveraging open-dictionary segmentation networks for continuous mapping. To summarize, our contributions are:
@@ II-A Continuous Semantic Mapping
Robots require advanced levels of scene understanding to plan, including knowledge of the geometry and semantic labels of objects and uncertainty associated with the objects to avoid failure due to mistaken object identity. Often, these approaches use task-dependent object designations as semantic labels [16] such as abstract topological information [21] or material classifications [22, 17]. Robotics research has focused on incorporating segmentation predictions into maps via semantic label fusion [23, 24]. Recent methods aim to quantify uncertainty through Bayesian inference [22, 17] by iteratively fusing semantic estimates projected onto a geometric map. Uncertainty quantification can be used by downstream planning algorithms to identify and circumvent potentially dangerous conditions [17], as well as to plan optimally informative trajectories in a field of robotics known as active perception [18, 19, 25].
Kernel-based inference schemes have had notable success [26, 4] in probabilistic semantic mapping. Bayesian Kernel Inference (BKI), proposed by [20], approximate the spatial influence of points at model selection through the usage of a kernel. BKI is an approximation of Gaussian Processes, which are effective at continuous mapping yet suffer from a cubic computational complexity [14, 27, 28]. Effectively, the kernel defines the shape or distribution of a point, deemed the extended likelihood, and can be applied to create an efficient, closed-form Bayesian update with more complete maps and quantifiable uncertainty [15]. While BKI has been applied effectively to semantic mapping [16, 4, 29], semantic maps are inherently limited to a closed set of pre-specified categories. In contrast, we propose to extend the literature of continuous mapping to the latent space of neural networks through a Gaussian likelihood and conjugate pair, allowing for open-vocabulary inference within the latent space of VL models, without any loss in performance.
Figure 2: This figure demonstrates the overall pipeline of LatentBKI. (a) The input to LatentBKI is 3D points, which can be from LiDAR, RGB-D, or any exteroceptive sensor with 3D input. (b) Points are then processed by an off-the-shelf neural network, which encodes each point into a latent space. (c) By adopting a Gaussian likelihood over the point-wise features, we perform closed-form Bayesian inference on a voxel map where each voxel contains parameters modeling the conjugate prior of the multivariate Gaussian distribution. Additionally, instead of only considering points which fall within a voxel, we consider nearby points weighted through a kernel function. (d) The posterior predictive distribution of each voxel in latent space can then be decoded using the decoder of the neural network, enabling the computation of open-dictionary segmentation predictions with expectation and uncertainty.
@@ II-B Vision Language Mapping
Rapidly improving Large Language Models (LLMs) demonstrating remarkable generalizable capabilities have motivated the advent of vision-language models (VLM) with shared latent space for both images and texts [7, 30, 31]. The pioneering method CLIP successfully represents visual and textual information in the same embedding dimension through contrastive learning [6]. Trained on a large dataset of image-text pairs, CLIP learns to embed features from visual or textual information in a shared feature space, where similarity is measured by a cosine similarity function [6].
Based on the success of VLMs and their great zero-shot performance, recent robotics research has focused on open-dictionary mapping which operates in the latent space of VLMs and can create segmentation predictions from language descriptors [8]. Approaches like LM-Nav [32], CoW [9], NLMap [12] and VLMap [10] have fused VLMs to enable robots to understand and navigate new environments. One common approach in literature to open-dictionary mapping is a volumetric averaging technique [11, 10, 13] where images are processed through a language-driven semantic segmentation network such as LSeg [8], producing 3D points paired with neural features. Points are then incrementally fused within a volumetric map structure as a moving average of the features of points falling within each map cell. At inference time, the latent expectation of each voxel can then be decoded into per-category scores given the language embeddings of a set of categories, thereby enabling open-dictionary queries. While successful, this approach loses the ability to quantify uncertainty and fill in gaps in the map from probabilistic continuous mapping, as discussed previously. Therefore, we propose to extend closed-dictionary continuous semantic mapping to open-dictionary VL maps to obtain quantifiable uncertainty and spatial smoothing.
@@ III-A Latent Mapping Representation
Our map representation consists of a voxel map with voxels * located at position $\bm{x}_{*}$ At each time-step our map is provided a set of points $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{3}$ is the position of point $i$ and $\bm{y}_{i}\in\mathbb{R}^{C}$ is the corresponding latent feature of point $i$. From these points, our goal is to probabilistically update the latent parameters of each voxel, $\bm{y}_{*}$, to obtain the posterior.
@@
In order to accomplish this goal, we first define a Gaussian likelihood over the feature space, such that: $p(\bm{y}_{i}|\bm{\mu}_{i},\Sigma_{i})=\mathcal{N}(\bm{y}_{i};\bm{\mu}_{i},\Sigma_{i})$. Since the features originate from a neural network, the likelihood defines an unknown expectation and variance which the point’s features are sampled from. Similarly, we can define the points observed within a voxel * according to the same likelihood distribution. From the likelihood, we can write an expression for the posterior of the latent parameters $\theta_{*}=\{\bm{\mu}_{*},\Sigma_{*}\}$ of voxel * using Bayes’ rule as:
$p(\theta_{*}\mid\bm{x}_{*},\mathcal{D})\propto p(\mathcal{D}\mid\theta_{*},\bm{x}_{*})p(\theta_{*}\mid\bm{x}_{*}).$ (3)
@@
In order to model the distribution over the parameters $\theta_{*}$ of the voxel, which themselves define a multivariate Gaussian distribution, we adopt the conjugate prior of the multivariate Gaussian distribution, the normal-inverse Wishart distribution. The normal-inverse Wishart distribution defines a distribution over the multivariate Gaussian with unknown mean $\bm{\mu}_{*}$ and covariance $\Sigma_{*}$ through parameters $\bm{\mu}^{\prime}_{*},\Psi_{*},\lambda_{*},v_{*}$ as:
@@
$p(\bm{\mu}_{*},\Sigma_{*})=\mathcal{N}\left(\bm{\mu}_{*};\bm{\mu}^{\prime}_{*},\frac{\Sigma_{*}}{\lambda_{*}}\right)\cdot\mathcal{W}^{-1}(\Sigma;\Psi_{*},v_{*}),$ (4)
@@
where the hyper-parameters represent the prior expectation of the mean ($\bm{\mu}^{\prime}_{*}$), the expectation of the covariance ($\Psi_{*}$), and the confidence in the mean and covariance estimates ($\lambda_{*},v_{*}$). In our case, $\lambda_{*}$ and $v_{*}$ are equal and correspond to weighted counts of the total observations.
@@
Although the conjugate prior provides a closed-form solution for updating the multivariate normal distribution parameters for each voxel, it does not consider the spatial locations of points. Intuitively, points that are closer to the centroid of the voxel should have a higher influence, while points that are further from the voxel centroid should have a lower influence. Additionally, only updating the voxels which points fall into can lead to sparse maps, as previously noted. Therefore, we adopt the solution of [20], which defines an extended likelihood distribution that considers the spatial relationship of points to voxels through the use of a kernel function $k(\bm{x}_{i},\bm{x}_{*})$ as:
$p(\bm{y}_{i}|\bm{x}_{i},\theta_{*},\bm{x}_{*})\propto p(\bm{y}_{i}|\theta_{*})^{k(\bm{x}_{i},\bm{x}_{*})}.$ (5)
@@
The only requirements when defining the extended likelihood are that $\color[rgb]{0,0,0}k(\bm{x},\bm{x})=1\ \forall x$ and $\color[rgb]{0,0,0}k(\bm{x},\bm{x}_{*})\in[0,1]\,\forall\,(\bm{x},\bm{x}_{*}).$ Applied to the previously defined Gaussian likelihood, the extended likelihood can be written as:
$p(\bm{y}_{i}|\theta_{*},\bm{x}_{i},\bm{x}_{*})=\mathcal{N}\left(\bm{y}_{i};\bm{\mu}_{*},\frac{\Sigma_{*}}{k(\bm{x}_{i},\bm{x}_{*})}\right).$ (6)
Following Semantic BKI, we use a symmetric sparse kernel [33] with kernel length $l=0.5$ for direct comparison, where $d$ is the Euclidean distance between the two points:
$k(d)=$
$\begin{cases}[\frac{1}{3}(2+\cos(2\pi\frac{d}{l})(1-\frac{d}{l})+\frac{1}{2\pi}\sin(2\pi\frac{d}{l}))],&\text{if }d<l\\ 0.&\text{else}\end{cases}$ (7)
Substituting the extended likelihood into (3), we can now define a spatial update over the voxel parameters as:
$p(\theta_{*}|\bm{x}_{*},\bm{y}_{1:N},\bm{x}_{1:N})\propto\left[\prod_{i=1}^{N}p(\bm{y}_{i}|\theta_{*},\bm{x}_{i},\bm{x}_{*})\right]p(\theta_{*}|\bm{x}_{*}).$ (8)
@@ III-B Latent Mapping Update
First, we initialize the confidence over the mean and covariance of each voxel to a non-informative value of $\lambda_{*}\approx 0$. As points are observed, the value of $\lambda_{*}$ increases, indicating more confidence in the expected mean and covariance of the voxel. At time-step $t$, voxels are parameterized by prior $\bm{\mu}_{*}^{t-1}$, $\Psi_{*}^{t-1}$ and confidence $\lambda_{*}^{t-1}$, with input points $\mathcal{D}$.
The influence of the new observations is calculated by: $\bar{k}_{*}=\sum_{i=1}^{N}k(\bm{x}_{*},\bm{x}_{i})$, where the kernel function measures the influence of point $i$ over voxel *. The confidence in the mean and covariance is updated: $\lambda_{*}^{t}=\lambda_{*}^{t-1}+\bar{k}_{*}$. Input observations are then used to compute the new mean as a running average: $\bar{\bm{y}}_{*}=\sum_{i=1}^{N}\frac{k(\bm{x}_{*},\bm{x}_{i})}{\bar{k}_{*}}\bm{y}_{i}$, $\bm{\mu}_{*}^{t}=\frac{\lambda_{*}^{t-1}\bm{\mu}_{*}^{t-1}+\bar{\bm{y}}_{*}\bar{k}_{*}}{\lambda_{*}^{t}}$.
@@
As new points are obtained, the above process is repeated to update the mean, covariance, and confidence level.